# Regulating Sommerfeld resonances for multi-state systems and higher partial waves

### v1

@@ 3.5 Characterizing the effects from short-range physics
If we wish to explicitly separate out the short-distance behavior controlled by $f_{s,\ell}$, it is easiest to work in the convention where $V_{\text{long}}(r)=0$ for $r<a$. Then as discussed above, $\kappa_{\ell}^{-1}=\left(f_{s,\ell}^{-1}\tilde{P}^{2\ell}-\tilde{P}\alpha_{\tilde{G}_{\ell}}(a)\right)(\tilde{P}^{\dagger})^{-2\ell}$. It is notationally convenient to define $\bar{f}_{s,\ell}\equiv(\tilde{P}^{\dagger}\tilde{P}^{-1})^{2\ell}f_{s,\ell}$, $\bar{Z}\equiv\tilde{P}\alpha_{\tilde{G}_{\ell}}(a)(\tilde{P}^{\dagger})^{-2\ell}$, in which case $\kappa_{\ell}^{-1}=\bar{f}_{s,\ell}^{-1}-\bar{Z}$. Then Eq. 144 takes the form:
$(\sigma_{i,\text{ann}}v_{\text{rel}})_{\ell}$ $=c_{i}\frac{2\pi i}{\mu}(2\ell+1)\left[\Sigma_{0,\ell}^{\dagger}\left[(f_{s,\ell}^{\dagger})^{-1}-\bar{Z}^{\dagger}+i\Sigma_{0,\ell}P^{2}\Sigma_{0,\ell}^{\dagger}\right]^{-1}\left(\bar{f}_{s,\ell}^{-1}-(\bar{f}_{s,\ell}^{\dagger})^{-1}+\bar{Z}^{\dagger}-\bar{Z}\right)\right.$
$\left.\left[f_{s,\ell}^{-1}-\bar{Z}-i\Sigma_{0,\ell}P^{2}\Sigma_{0,\ell}^{\dagger}\right]^{-1}\Sigma_{0,\ell}\right]_{ii}$
$=c_{i}\frac{2\pi i}{\mu}(2\ell+1)\left[\bar{\Sigma}_{\ell}^{\dagger}\left(\bar{f}_{s,\ell}^{\dagger}-\bar{f}_{s,\ell}+\bar{f}_{s,\ell}^{\dagger}(\bar{Z}^{\dagger}-\bar{Z})\bar{f}_{s,\ell}\right)\bar{\Sigma}_{\ell}\right]_{ii},$
$\bar{\Sigma}_{\ell}$ $=\left[1-(\bar{Z}+i\Sigma_{0,\ell}P^{2}\Sigma_{0,\ell}^{\dagger})\bar{f}_{s,\ell}\right]^{-1}\Sigma_{0,\ell}.$ (146)
However, because $\bar{Z}$ is in general non-Hermitian, the numerator in this case has a more complicated form than in Eq. 145. To maintain the form of the cross section given in Eq. 145 but make $\tilde{\kappa}_{\ell}$ independent of the long-range physics, we could choose a Hermitian $\Delta_{\ell}$ to absorb the dependence on the long-range potential, i.e. $\Delta_{\ell}=\frac{1}{2}(\bar{Z}+\bar{Z}^{\dagger})$ (the Hermitian part of $\bar{Z}$). Then $\tilde{\kappa}_{\ell}^{-1}=\bar{f}_{s,\ell}^{-1}-\frac{1}{2}(\bar{Z}-\bar{Z}^{\dagger})$. It turns out that the anti-Hermitian part of $\bar{Z}$ holds no information on the long-range potential, so this choice moves all the terms dependent on the long-range potential into $\Delta_{\ell}$.
More specifically, $\bar{Z}^{\dagger}-\bar{Z}=i(\tilde{P}^{-2\ell}(\tilde{P}^{\dagger})^{2(\ell+1)}+\tilde{P}^{2(\ell+1)}(\tilde{P}^{\dagger})^{-2\ell})$. To prove this, note that $\tilde{G}_{ni}(a)$ is real by construction (it was defined as a real wavefunction), and we have
@@
Since $s_{\ell}(p_{n}a)p_{n}^{-(\ell+1)}$, $c_{\ell}(p_{n}a)p_{n}^{\ell}$, and $G_{ni}(a)$ are all real, it follows that $\tilde{P}^{2\ell+1}(\alpha_{\tilde{G}_{\ell}}(a)+i\tilde{P}^{2\ell+1})$ must also be real. Furthermore, we show in App. B that $\tilde{P}^{2\ell+1}\alpha_{\tilde{G}_{\ell}}(0)=\tilde{P}^{2\ell+1}\alpha_{\tilde{G}_{\ell}}(a)$ must also be symmetric, by direct construction. Thus $\tilde{P}^{2\ell+1}(\alpha_{\tilde{G}_{\ell}}(a)+i\tilde{P}^{2\ell+1})$ is Hermitian, and so is $\tilde{P}^{-2\ell}\tilde{P}^{2\ell+1}(\alpha_{\tilde{G}_{\ell}}(a)+i\tilde{P}^{2\ell+1})(\tilde{P}^{\dagger})^{-2\ell}=\bar{Z}+i\tilde{P}^{2(\ell+1)}(\tilde{P}^{\dagger})^{-2\ell}$. The required relationship follows directly.
We can now interpret the “numerator” term in Eq. 3.5. $\bar{Z}^{\dagger}-\bar{Z}$ is a diagonal matrix with the $n$th diagonal component being $2ip_{n}$ for $n\leq M$, and zero for $n>M$. When we write out the non-Hermitian part of the short-range scattering amplitude, according to the optical theorem it will include a contribution corresponding to the cross section for scattering into kinematically allowed states (i.e. $n\leq M$), but this contribution is not considered part of our inclusive annihilation cross section and must be subtracted. This subtraction is implemented by the $\bar{f}_{s,\ell}^{\dagger}(\bar{Z}^{\dagger}-\bar{Z})\bar{f}_{s,\ell}$ term in Eq. 3.5. We discuss this issue more explicitly in App. C.
@@ 3.6.1 Case with no long-range potential
Let us first consider the case where we turn off the long-range potential completely, setting $V_{\text{long}}(r)=0$ (both inside and outside $r=a$). Then we have $N=M$, $Q_{\ell}=\Sigma_{0,\ell}=1$, and $f_{s.\ell}=\hat{f}_{s,\ell}$. The wavefunctions for the long-range potential are just the free-particle wavefunctions, and $\tilde{G}_{ni}(r)=\delta_{ni}p_{n}^{\ell}c_{\ell}(p_{n}r)=\delta_{ni}p_{n}^{\ell+1/2}(g_{n}(r)-if_{n}(r))$ (note that in this case $p_{n}$ is always real as $M=N$). Thus we can read off $\alpha_{\tilde{G}_{\ell}}(a)=-iP^{2\ell+1}$, and

### v3

@@ 3.5 Characterizing the effects from short-range physics
If we wish to write the cross section directly in terms of $\bar{f}_{s,\ell}$ rather than $\kappa_{\ell}$, we can expand Eq. 146 as:
$(\sigma_{i,\text{ann}}v_{\text{rel}})_{\ell}$ $=c_{i}\frac{2\pi i}{\mu}(2\ell+1)\left[\Sigma_{0,\ell}^{\dagger}\left[(\bar{f}_{s,\ell}^{\dagger})^{-1}-\bar{Z}_{\ell}^{\dagger}+i\Sigma_{0,\ell}P^{2}\Sigma_{0,\ell}^{\dagger}\right]^{-1}\left(\bar{f}_{s,\ell}^{-1}-(\bar{f}_{s,\ell}^{\dagger})^{-1}+\bar{Z}_{\ell}^{\dagger}-\bar{Z}_{\ell}\right)\right$.
$\left.\left[\bar{f}_{s,\ell}^{-1}-\bar{Z}_{\ell}-i\Sigma_{0,\ell}P^{2}\Sigma_{0,\ell}^{\dagger}\right]^{-1}\Sigma_{0,\ell}\right]_{ii}$
$=c_{i}\frac{2\pi i}{\mu}(2\ell+1)\left[\bar{\Sigma}_{\ell}^{\dagger}\left(\bar{f}_{s,\ell}^{\dagger}-\bar{f}_{s,\ell}+\bar{f}_{s,\ell}^{\dagger}(\bar{Z}_{\ell}^{\dagger}-\bar{Z}_{\ell})\bar{f}_{s,\ell}\right)\bar{\Sigma}_{\ell}\right]_{ii},$
$\bar{\Sigma}_{\ell}$ $=\left[1-(\bar{Z}_{\ell}+i\Sigma_{0,\ell}P^{2}\Sigma_{0,\ell}^{\dagger})\bar{f}_{s,\ell}\right]^{-1}\Sigma_{0,\ell}.$ (148)
However, because $\bar{Z}_{\ell}$ is in general non-Hermitian, the numerator in this case has a more complicated form than in Eq. 147. It is easiest to interpret the additional $\bar{f}_{s,\ell}^{\dagger}(\bar{Z}_{\ell}^{\dagger}-\bar{Z}_{\ell})\bar{f}_{s,\ell}$ numerator term if we use the convention where we set $V_{\text{long}}(r)=0$ for $r<a$, so that $\bar{f}_{s,\ell}=(\tilde{P}^{\dagger}\tilde{P}^{-1})^{2\ell}f_{s,\ell}$ contains all the short-distance physics including scattering in the potential at $r<a$, and $\bar{Z}_{\ell}=\tilde{P}\alpha_{\tilde{G}_{\ell}}(a)(\tilde{P}^{\dagger})^{-2\ell}$.
In this convention, $\bar{Z}_{\ell}^{\dagger}-\bar{Z}_{\ell}=i(\tilde{P}^{-2\ell}(\tilde{P}^{\dagger})^{2(\ell+1)}+\tilde{P}^{2(\ell+1)}(\tilde{P}^{\dagger})^{-2\ell})$. To prove this, note that $\tilde{G}_{ni}(a)$ is real by construction (it was defined as a real wavefunction), and we have
@@
Since $s_{\ell}(p_{n}a)p_{n}^{-(\ell+1)}$, $c_{\ell}(p_{n}a)p_{n}^{\ell}$, and $G_{ni}(a)$ are all real, it follows that $\tilde{P}^{2\ell+1}(\alpha_{\tilde{G}_{\ell}}(a)+i\tilde{P}^{2\ell+1})$ must also be real. Furthermore, we show in App. B that $\tilde{P}^{2\ell+1}\alpha_{\tilde{G}_{\ell}}(0)=\tilde{P}^{2\ell+1}\alpha_{\tilde{G}_{\ell}}(a)$ must also be symmetric, by direct construction. Thus $\tilde{P}^{2\ell+1}(\alpha_{\tilde{G}_{\ell}}(a)+i\tilde{P}^{2\ell+1})$ is Hermitian, and so is $\tilde{P}^{-2\ell}\tilde{P}^{2\ell+1}(\alpha_{\tilde{G}_{\ell}}(a)+i\tilde{P}^{2\ell+1})(\tilde{P}^{\dagger})^{-2\ell}=\bar{Z}_{\ell}+i\tilde{P}^{2(\ell+1)}(\tilde{P}^{\dagger})^{-2\ell}$. The required relationship follows directly.
Consequently, in this convention $\bar{Z}_{\ell}^{\dagger}-\bar{Z}_{\ell}$ is a diagonal matrix with the $n$th diagonal component being $2ip_{n}$ for $n\leq M$, and zero for $n>M$. When we write out the non-Hermitian part of the short-range scattering amplitude, according to the optical theorem it will include a contribution corresponding to the cross section for scattering into kinematically allowed states (i.e. $n\leq M$), but this contribution is not considered part of our inclusive annihilation cross section and must be subtracted. This subtraction is implemented by the $\bar{f}_{s,\ell}^{\dagger}(\bar{Z}_{\ell}^{\dagger}-\bar{Z}_{\ell})\bar{f}_{s,\ell}$ term in Eq. 3.5. We discuss this issue more explicitly in App. D.
@@ 3.6.1 Case with no long-range potential
Let us first consider the case where we turn off the long-range potential completely, setting $V_{\text{long}}(r)=0$ (both inside and outside $r=a$). In this case the convention choice for $r<a$ is trivial. Then we have $N=M$, $Q_{\ell}=\Sigma_{0,\ell}=1$, and $f_{s.\ell}=\bar{f}_{s,\ell}$. The wavefunctions for the long-range potential are just the free-particle wavefunctions, and $\tilde{G}_{ni}(r)=\delta_{ni}p_{n}^{\ell}c_{\ell}(p_{n}r)=\delta_{ni}p_{n}^{\ell+1/2}(g_{n}(r)-if_{n}(r))$ (note that in this case $p_{n}$ is always real as $M=N$). Thus we can read off $\alpha_{\tilde{G}_{\ell}}(a)=-iP^{2\ell+1}$, and